OpenStudy (anonymous):

find limit as X reaches infinity 5x^2 -8x-13

OpenStudy (anonymous):

\(\infty\).

OpenStudy (anonymous):

My infinity looks much cooler than yours kushashwa23 :P

OpenStudy (anonymous):

but if i take it's derivative and then try to find the answer, shouldn't that work?

OpenStudy (anonymous):

wait : \[∞\]

OpenStudy (anonymous):

\[\color{red}{∞}\]

OpenStudy (anonymous):

so 4/5 should be the answer right?

OpenStudy (anonymous):

No! I think you're talking about L'Hopital's Rule, which is only applicable if you have something of the form \(\frac{0}{0}\) or \(\frac{\infty}{\infty}\).

OpenStudy (anonymous):

anwarA now mine is cooler one :P \[\color{ blue}{∞}\]

OpenStudy (anonymous):

I agree, and you deserve a medal for that!

OpenStudy (anonymous):

Call me Anwar.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

but if i try to plot the graph, it has a zero gradient at x=8, so the limit would be where x= 4/5

OpenStudy (anonymous):

and putting x=4/5 in the actual equation i can get the limit right?

OpenStudy (anonymous):

ans will be \[\infty\]

OpenStudy (anonymous):

read my suggestion and let me know if my approach is wrong or not

OpenStudy (anonymous):

I don't think I get your question. See the graph here http://bit.ly/ncbWjB. You can see that the function goes to \(\infty\) as you take large values of \(x\).

OpenStudy (anonymous):

okay=)